OpenStudy (anonymous):

Solve by substitution. y = 2x + 3 5x - y = -3

OpenStudy (anonymous):

plug the first equation into the 2nd equation and get 5x - (2x+3) = -3 solve for x 5x - 2x - 3 = -3 3x = 0 x = 0 y = 2(0)+3 y=3

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

yea